# Surveying the landscape of optically addressable spin qubits for quantum information and sensing technology

### v1

@@ 2 Overview of Candidate Spin Qubit Materials
Beyond vacancy systems, several metal ion centres have demonstrated interesting spin-optical properties. By far the best however is Cr4+ in 4H-SiC. As a S=1 species, this material exhibits >1 second, and and = 81 s and 317 ns, respectively, at 15 K[55]. Importantly, it also demonstrates a 79 contrast, marking it as a system with some of the highest optical readout fidelity. The spin-optical properties are markedly impaired in GaN host, which demonstrates 27x broader emission linewidths due to interactions of Cr4+-centres with the surrounding spin bath[54].
p- and d-block Molecular FOAMs
@@
Molecular systems are becoming increasingly popular and offer an enticing opportunity to develop chemically tuneable quantum materials catered to different applications[98, 99, 100, 101]. Ground-up synthesis enables the incorporation of particular functionalities such as stable radicals[102, 103], modulation of triplet/singlet yields, enrichment with low or zero nuclear magnetic moments such as deuterium, oxygen and sulfur, or the targeted inclusion of nuclear spin-active elements such as nitrogen, phosphorous and several transition metals and lanthanides (vide infra). A synthetic approach also enables changes to the host matrix[73], spin concentration, defect orientation, and material processing approaches which are limited for defect-based systems. For example, single-molecule ODMR spectroscopy of pentacene molecules in a p-terphenyl matrix (Pc:PTP) has already been demonstrated at cryogenic temperature in a series of remarkable works by Wratchup and colleagues[104, 105, 106]. Only recently have pulsed experiments been performed to reveal highly competitive contrast and spin coherence properties at RT. The Tx-Ty spin transition of a 0.1 crystal Pc:PTP demonstrates 23 s, 2.7 s and 500 ns by pulsed ODMR spectroscopy[69, 107]. Investigations using the more strongly spin polarised Tx-Tz transition in both crystals and 100 nm-thin films at 0.01 and 0.1, respectively, reveal similar spin dynamics and also suggest an ability to modulate Pc:PTP’s spin properties according to sample thickness and spin concentration[70].
@@
Further organic systems have demonstrated potential as FOAMs, though to our knowledge pulsed optically detected experiments have yet to be performed. For example, the room temperature steady-state ODMR contrast of a 1 crystal of pentacene-doped picene has been measured at 15, representing a potentially significant improvement over the PTP matrix[108]. Steady-state ODMR signals have also previously been reported at 2 K for (perdeutero)tetracene, 1,2-benzathracene, 1,2,3,4-dibenzanthracene[109] and dinaphtho-(2’,3’:1,2);(2",3":6,7)-pyrene[110]. Interestingly, work by Corvaja, Pasimeni and Giometti et al., have shown that even highly spin-dense charge-transfer (CT) co-crystals can exhibit bright RT ODMR signals. Co-crystals comprised of donors such as biphenyl, fluorene, phenazine and acceptors such as 7,7’:8,8’-tetracyanoquinodimethane (TCNQ) and 1,2,3,4-tetrafluoro-TCNQ (F4TCNQ) have been studied to elucidate their triplet state dynamics[111, 112]. These materials exhibit narrow resonance lines due to intermolecular site hopping. The resonances for each site can become resolved at low temperatures where hopping is not thermodynamically favoured[111]. Their high spin densities (50) are highly advantageous for quantum sensing where the a.c. sensitivity is proportional to the , though this is often concurrent with less robust quantum spin properties compared with dilute materials such as Pc:PTP. Moreover, these materials are also promising hosts for exotic spin behaviours like singlet fission and triplet-triplet annihilation[113]. For example, at RT the triplet states of Phenazine:TCNQ are predominately formed by singlet fission. Despite high spin densities, experiments using transient nutation ESR spectroscopy have been used to estimate and times of 1 s and 600 ns, respectively[114]. More recently, several ground state radical and diradical materials with an optical readout capacity have also been demonstrated[115, 71], as have materials with quintet states[116, 117, 118].
FOAMs consisting of d-block have only recently been demonstrated with Cr4+. The Cr4+ spin-centre was realised with several tolyl-based ligand systems where modifications to the ligand structure and corresponding ligand field strength can be used to control optical excitation frequency and the ZFS of the Cr4+ S=1 ground state[119]. To reduce dipole coupling, the spin-active moiety is diluted in a matrix of the S=0 isostructural tin analogue with crystals grown from hexane solutions[72]. Optical spin polarisation is then achieved by exciting molecular spins between the S=1 ground state and S=0 excited state, which fluorescently decays within a few microseconds to favourably populate the T± states. Initial pulsed ODMR experiments were performed on Cr4+(o-tolyl)4 spins which benefit from an E=0 triplet state and a relatively long of 0.22 ms at 5 K. This permits several optical cycles to build spin polarisation and enhance the optical contrast, enabling the measurement of of 640 ns. The coherent properties can be significantly improved by using a non-isostructural tin(4-fluoro-2-methylphenyl)4 host matrix[120]. Here, the ZFS is significantly increased giving rise to clock transitions similar to those realised by the ZEFOZ method. As a result, and are increased to 1.210.02 ms and 10.60.2 s at 5 K, respectively.
@@
Recently, even fluorescent proteins have been demonstrated as a viable spin qubit media in vivo and in solution at room temperature. Feder et al., used ODMR spectroscopy to demonstrate that at 80 K the X-Z and Y-Z triplet transitions of enhanced yellow fluorescent protein (EYFP) demonstrated a 44 and 32 ODMR contrast, respectively[97]. This corresponds with zero-applied field (estimated from spin polarisation decay) and of 141 and 1.5 s. Using Carr-Purcell-Meiboom-Gill (CPMG) dynamical decoupling the effective decoherence time () reached 16 s. The authors largely circumvent overhead limitations from long triplet lifetimes (ms) using an additional near-infrared pulse to induce T1 to T2 transitions thereby inducing reverse intersystem crossing and delayed fluorescence from S1. This interesting methodology could be suitable for other spin systems with quasi-resonant S1 and T2 electronic states, such as pentacene, to increase their sensitivity by increasing measurement repetition rates.
f-block molecular FOAMs
@@
Lastly, there has been significant interest in d- and f-block molecular systems. Here, we make a distinction between molecular systems, where the ion is doped into a molecular lattice, and trapped ion systems which can give rise to extremely long coherence times, but use radically different device architectures and are not subject to qubit-quality improvements through molecular engineering[121, 122, 123]. Like organic materials, d- and f-block molecular systems benefit from diverse chemistry and processing methods that can in principle reduce decoherence, enable built-in error correction, and lead to scalable qubit media. Moreover, these heavy elements often exhibit strong spin-orbit coupling leading to high magnetic anisotropy that can protect spin states from small magnetic fluctuations[124, 125]. However, as a result of these effects and unlike the candidates previously discussed, the optical transitions in question often correspond to nuclear spin transitions. Examples of optically active materials benefit from narrow and stable spin-dependent emission profiles at near-infrared frequencies making them potentially compatible with conventional telecom optical fibres. Colour centres for which optical addressability has been established include Cr4+[72], Eu3+-[126], Pr3+-[92], Er3+-[89], Yb3+-[127], and Sm3+:Y2SiO5 [128]. These lanthanide-based centre benefit from highly shielded f-orbital electrons leading to relatively long coherence times at low temperatures[129], as well as strong hyperfine coupling (tens of MHz) with I=5/2 (Eu, Pr, 173Yb), 1/2 (171Yb), 7/2 (149Sm, 167Er) nuclei combined with strong spin-orbit coupling leading to discreet optically addressable spin-sublevels. Y2SiO5 has been favored due to the ability to grow large crystals with excellent optical properties by the Czochralski method[130]. However, as the only naturally occurring isotope, 89Y harbours an I=1/2 nuclear spin that ultimately limits decoherence times. To reduce the impact of the spin bath and inhomogeneity of applied-magnetic fields, various decoupling techniques have emerged[131, 132, 87]. Perhaps the most successful is the so-called zero first-order Zeeman (ZEFOZ) technique, whereby a magnetic field is applied such that the magnetic-field dependence of the spin-transition frequency is very close to zero [133]. In this "clock transition" regime, spins are first-order insensitive to small fluctuations in local magnetic fields. Using this technique with Eu3+:Y2SiO5, Zhong et al., demonstrated it is possible to acquire >>100 ms at 2 K, where spin-phonon coupling is negligible. Remarkably, combined with dynamic decoupling methods was measured to be 37060 mins, reaching a critical milestone whereby the distance-dependent decoherence becomes less for spin-transport than it is in light-transport of quantum states[134].
@@
Interestingly, the larger magnetic moment of Pr3+ can give rise to "frozen core"-type behaviour, whereby local Y-spins become dephased from the bulk crystal and hence exhibit reduced dephasing influence on the Pr3+ spins[131]. Equall et al., measured homogenous field-dependent and crystal structure site-dependent linewidths between 2.5 and 0.85 kHz, and a corresponding as high as 377 s at 1.4 K[92]. Combined with the ZEFOZ method, the can reach 82 ms at 1.5 K[131] and with further dynamical decoupling up to 1 min can be achieved, approaching the population lifetime limit[135]. Coherence times can be further enhanced by using a host matrix whereby the principle host ion (e.g., Y) has a more closely matched ionic radius to the dopant, leading to reduced crystallographic distortions. For example, in a Pr3+:La2(WO4)3 system a of 1587 ms has been measured at 4 K[93].
@@
Spin coherence times of Er3+:Y2SiO5 have so far been shorter than the best performing Eu3+ materials with measured at 1.30.01 seconds at 1.4 K, despite exhibiting the frozen core effect[89]. However, this was achieved without using the ZEFOZ method and therefore these times can likely be significantly extended. Er3+ has also been studied in a Y2O3 host with reaching 140 s at 1.8 K and with an applied field of 4 T[88]. Decoherence was dominated by phonon-driven dipole-dipole interactions and the nuclear spin bath at high fields, similar to Er3+:KTiOPO4 where was measured at 200 s under similar conditions[90].
@@
171Yb3+ holds a unique position amongst the lanthanide ions discussed so far due to its S=1/2 and I=1/2 electron spin and hyperfine structure resulting in a simple 4-level system. Of the 171Yb3+-doped materials[95], 171Yb3+:Y2SiO5 appears to exhibit the longest spin coherence times. This material was initially studied by X-band EPR spectroscopy and presented with an electron of 5 seconds at 2.5 K and 90 mT, which quickly increases above 4 K due to Raman relaxation where is ultimately limited by [94]. At 2.5 K, was optimised at 1 T to 73 s and improved further by dynamical decoupling until reached 550 s. In the same experiments, the authors record nuclear and at 4.5 K of 4 and 0.35 ms, respectively. Using an optical approach, Ortu et al., improved the coherent properties of 171Yb3+:Y2SiO5 by employing the ZEFOZ method such that the electron remains above 100 s at 5.6 K and the nuclear extend 1 ms[127]. Using a different approach, Welinski et al., demonstrated that coherence can also be extended by first polarising host nuclear spins through spin diffusion. At 2 K, the authors first excite 171Yb spins before allowing them to equilibrate over a few seconds through spectral diffusion over the inhomogeneous linewidth. The result is an effective "hole burning" in the absorption spectrum of 171Yb3+:Y2SiO5 and up to 90 spin polarisation, thereby effectively generating mK spin temperatures. The result is an increase in the optical from 0.3 to 0.8 ms [136]. To our knowledge, pulsed optical decoherence studies have not been performed on Sm3+:Y2SiO5, however, its I=7/2 nucleus and strong hyperfine coupling may be useful for qudit systems, and it is predicted to be less sensitive to magnetic field fluctuations that Er3+[128, 137].
3 Guidelines for Optimising Performance of Spin-Qubit Materials

### v2

@@ 2.1.3 Inorganic Semiconductor Spin Centres
Beyond vacancy systems, several metal-ion centres have demonstrated interesting spin-optical properties. By far the best, however, is Cr4+ in 4H-SiC. As a S=1 species, this material exhibits >1 second, and and = 81 s and 317 ns, respectively, at 15 K [91]. Importantly, it also demonstrates a 79 contrast, marking it as a system with some of the highest optical readout fidelity. The spin-optical properties are markedly impaired in GaN host, which demonstrates 27x broader emission linewidths due to interactions of Cr4+-centres with the surrounding spin bath [41].
2.2 2D- and van der Waals FOAMs
Van der Waals (vdW) materials are made of 2D-atomically thin layers weakly bound together through van der Waals interactions. These include graphene, hexagonal boron nitride (hBN) and transition metal dichalcogenides (TMDs). Despite a range of remarkable materials having been identified as ODMR-active, the precise structure and even spin multiplicity of the spin species is often ambiguous. Nevertheless, like other defect spin-qubit systems, inhomogeneous broadening and spin-phonon interactions are the principal causes of spin decoherence.
The prototypical vdW material, graphene, was an early candidate for harbouring spin qubits; however, the metallic band structure of graphene is not conducive to forming luminescent spin centres. Instead, its isoelectronic analogue, hBN, has been the centre of remarkable advancement as a 2D-optically addressable platform at room temperature, building off its success as a single-photon source [103]. Magnetically-active spin centres in hBN were identified by EPR in the 1960s, with hyperfine structures giving clues to their identity as S=1/2 boron centres and carbon impurities [104, 105, 106]. However, the first optically addressable spin centres were only identified recently as triplets stemming from a negatively charged vacancy, likely originating from a missing boron atom, so-called -centres [107, 108, 109, 110] (Figure 6a). Similarly to NV-diamond centres, optical initialisation is achieved by photoexcitation with 532 nm light, leading to ISC into a metastable singlet, which spin-selectivity repopulates the fluorescent ms=0 state (TZ at zero-applied field, Figure 6b). The readout of the spin state occurs through 850 nm fluorescence with pulsed ODMR contrast of 1 at room temperature and almost 30 at 10 K, revealing a ZFS of |D|3.5GHz (Figure6c). Spin-relaxation and Hahn-echo measurements reveal a room-temperature at of 18 s and of 2 s [27]. These properties are particularly remarkable since both boron and nitrogen exhibit isotopically-abundant nuclear spins ( = 1, = 3, and = 3/2). Suppression of the nuclear spin bath using a second microwave source centred on a hyperfine line was found to improve and up to 25 s and 7.5 s, respectively. Work continues optimising the synthesis conditions for hBN to enhance the quantum properties of the defects. For instance, the contrast of -centres can be increased to 10 at room temperature by controlling the material purity and electron irradiation dose, though appears to remain phonon limited [111].
Due to hBN’s potential as a tunable single photon source, significant effort has been made to control carbon-defect structures through careful synthesis. CVD-grown hBN has been found to contain a plethora of optically-addressable carbon-based spin species with stable ZFLs in the visible range; though most have yet to be conclusively structurally characterised, or even the multiplicity conclusively determined (see Figure 6d for candidate structures and Figure 6e for the spin initialisation scheme). For example, three distinct so called ""-defects, likely stemming from the substitution of a boron or nitrogen atom by carbon ("D1", "D2", "D3", all likely S=1/2), demonstrate similar spin relaxation and coherence quantum properties (0.2 s) with a continuous wave (CW)-optical contrast up to 20 at 5 K [30] (Figure 6f). Their inhomogeneous linewidths are also limited by the nuclear spin bath (40-60 ns), where local spins also bestow an apparent zero-field splitting of between 8-10 MHz.
Though ensemble measurements remain challenging due to carbon-doped hBN’s broad emission profile, a recent report has shown that it is possible to survey over 400 defect centres in CVD-grown samples. Collectively known as ""-centres, careful measurements of individual spin-centres reveal distinctive triplet character with a ZFS of 15 MHz in contrast to most other accepted carbon-defects found in hBN [112]. The triplet identity was later unambiguously characterised and confirmed for a series of defects with significantly larger ZFS with |D|2 GHz and |E|50-80 MHz [29] (Figure 6g and h). The room-temperature was measured to be 20030 s, 1 s and 100 ns, competitive with -triplet centres. However, these properties are matched with a remarkable ODMR contrast of almost 50 (though it decreases with an applied magnetic field), giving rise to projected D.C. magnetic field sensitivity of up to 3 T Hz-1/2. A recent survey of similar defects demonstrated that under high microwave and optical pump powers, the optical contrast could reach as high as 90, though most exhibit <50 despite similar ZFS, and [113]. The authors explore the potential of this material as a medium for D.C. vectorial quantum magnetometry by taking advantage of the defects’ low symmetry and in-plane quantisation axis. Though yet-to-be experimentally verified, the authors hypothesise that applying a field bias along a different material axis corresponding to each of the three triplet states, Tx, Ty, and Tz, should result in subtle changes in contrast that can be related to particular Tn states to yield 3-vector components of the applied magnetic field.
More complex carbon species demonstrating similar coherent behaviour at room temperature have been explored using metal-organic vapour-phase epitaxy with carbon-based additives [114, 28]. The so-called ""-defect (potential identity ) is another potential S=1/2 species or, two weakly coupled S=1/2 centres, has recently been reported as being co-located with -centres following electron irradiation of hBN flakes [28]. While their and are similar to -species (13 s and 80 ns, respectively), it can be used to "sense" -centres through spin cross relaxation at an appropriate applied magnetic field. Due to its out-of-plane magnetic quantisation axis (versus the in-plane axis of -centres), it can also more effectively be used to image magnetic structure on underlying substrates with an estimated DC magnetic sensitivity of 1 T Hz-1/2.
2.3 Molecular FOAMs
2.3.1 p-block molecular spin centres
@@ 2.3 Molecular FOAMs
Molecular systems are becoming increasingly popular and offer an enticing opportunity to develop chemically tunable quantum materials catered to different applications [115, 116, 117, 118]. Ground-up synthesis enables the incorporation of particular functionalities such as stable radicals [119, 120], modulation of triplet/singlet yields, enrichment with low or zero nuclear magnetic moments such as deuterium, oxygen and sulfur, or the targeted inclusion of nuclear spin-active elements such as nitrogen, phosphorous, several transition metals, and lanthanides (vide infra). A synthetic approach also enables changes to the host matrix [19], spin concentration, defect orientation, and material processing approaches, which are limited for defect-based systems.
Single-molecule ODMR spectroscopy of pentacene molecules in a p-terphenyl matrix (Pc:PTP) has already been demonstrated at cryogenic temperatures in a series of remarkable works by Wrachtrup and colleagues [121, 122, 123]. Only recently have pulsed experiments been performed to reveal highly competitive contrast and spin coherence properties at room-temperature. The Tx-Ty spin transition of a 0.1 crystal Pc:PTP demonstrates 23 s, 2.7 s and 500 ns [13, 124]. Investigations using the more strongly spin polarised Tx-Tz transition in both crystals and 100 nm-thin films at 0.01 and 0.1, respectively, reveal similar spin dynamics and also suggests an ability to modulate Pc:PTP’s spin properties according to sample thickness and spin concentration [14]. These robust spin-optical properties can be maintained in Pc:PTP nanocrystals down to 200 nm with an optical contrast similar to equivalently sized NV-diamond nanoparticles [125]. When coated with a Pluronic F-127 polymer, these nanoparticles also display excellent biocompatibility with low toxicity over 72 hours according to in vitro explorations, paving the way for their use in cell assay techniques.
The potential of using synthetic chemistry to modulate the spin dynamics to improve spin-optical properties has also been explored. Substituting two carbon atoms on pentacene for nitrogen atoms resulted in 6,13-diazapentacene, which could be doped into a p-terphenyl host (DAP:PTP) by crystal growth and vapour deposition techniques [126, 127]. While maintaining a relatively long of 1.46 s and strong triplet spin polarisation, DAP:PTP exhibited an enhanced contrast of up to 40, surpassing the state of the art in NV-diamond platforms [127]. This improvement was attributed to the introduction of new low-energy non-bonding states that enhance vibrationally-induced spin-orbit coupling between the ground-state singlet and the metastable Tx state. It was also shown that direct solvent techniques can be used to grow ODMR-active DAP:PTP nanoparticles, which could be compatible with in vitro sensing experiments.
@@
Further organic systems have demonstrated potential as FOAMs, though, to our knowledge, pulsed optically detected experiments have yet to be performed. For example, the room temperature steady-state ODMR contrast of a 1 crystal of pentacene-doped picene has been measured at 15, representing a potentially significant improvement over the PTP matrix [130]. Steady-state ODMR signals have also previously been reported at 2 K for (perdeutero)tetracene, 1,2-benzathracene, 1,2,3,4-dibenzanthracene [131] and dinaphtho-(2’,3’:1,2);(2",3":6,7)-pyrene [132]. Interestingly, work by Corvaja, Pasimeni and Giometti et al., have shown that even highly spin-dense charge-transfer (CT) co-crystals can exhibit bright room-temperature ODMR signals. Co-crystals comprised of donors such as biphenyl, fluorene, phenazine and acceptors such as 7,7’:8,8’-tetracyanoquinodimethane (TCNQ) and 1,2,3,4-tetrafluoro-TCNQ (F4TCNQ) have been studied to elucidate their triplet state dynamics [133, 134]. These materials exhibit narrow resonance lines due to intermolecular site hopping of triplet excitons. The resonances for each site can become resolved at low temperatures where hopping is not thermodynamically favoured [133]. Their high spin densities (50) are highly advantageous for quantum sensing, where the a.c. sensitivity is proportional to the , though this is often concurrent with less robust quantum spin properties compared with dilute materials such as Pc:PTP. Moreover, these materials are also promising hosts for the study of exotic spin behaviours like singlet fission and triplet-triplet annihilation [135]. For example, at room-temperature, the triplet states of Phenazine:TCNQ are predominantly formed by singlet fission and using transient nutation ESR spectroscopy the authors estimate and times of 1 s and 600 ns, respectively, even at room temperature [136].
Several ground-state radical and diradical materials with an optical readout capacity have also been demonstrated [137, 15, 16], as have materials with quintet states [138, 139, 140]. Diradical materials have the advantage of exhibiting a ground-state triplet, making their optical addressability protocol identical to NV-diamond. Optical readout is enabled by employing luminescent radical moieties based on tris(2,4,6-trichlorophenyl)methyl (TTM) [141], where the radical is sterically isolated on a carbon atom with only weak hyperfine coupling to neighbouring 13C and hydrogen atoms. As a result, these materials can benefit from relatively narrow resonance linewidths, long s, and can be embedded in frozen solutions or optically transparent PMMA foils through simple solvent-based methods [129].
@@
Recently, even fluorescent proteins have been demonstrated as a viable spin qubit medium in vivo and in solution at room temperature. Feder et al., used ODMR spectroscopy to demonstrate that at 80 K the X-Z and Y-Z triplet transitions of enhanced yellow fluorescent protein (EYFP) demonstrated a 44 and 32 ODMR contrast, respectively [56]. This corresponds with a zero-applied field of 141 s (estimated from spin polarisation decay) and of 1.5 s. Using Carr-Purcell-Meiboom-Gill (CPMG) dynamical decoupling the effective decoherence time () reached 16 s. The authors largely circumvent overhead limitations from long triplet lifetimes (ms) using an additional near-infrared pulse to induce T1 to T2 transitions to induce reverse ISC and delayed fluorescence from S1. This interesting methodology could be suitable for other spin systems with quasi-resonant S1 and T2 electronic states, such as pentacene, to increase their sensitivity by increasing measurement repetition rates.
2.3.2 d-block molecular spin centres
Molecular FOAMs with d-block elements have only recently been demonstrated with Cr4+. The Cr4+ spin-centre was realised with several tolyl-based ligand systems where modifications to the ligand structure and corresponding ligand field strength can be used to control optical excitation frequency and the ZFS of the Cr4+ S=1 ground state [142]. To reduce intermolecular dipole coupling, the spin-active moiety is diluted in a matrix with the S=0 isostructural tin analogue in crystals grown from hexane solutions [18]. Optical spin polarisation is then achieved by exciting molecular spins from the S=1 state to the S=0 manifold, which fluorescently decays within a few microseconds to favourably populate the T± states. Initial pulsed ODMR experiments were performed on Cr4+(o-tolyl)4 spins, which benefit from an E=0 triplet and a relatively long of 0.22 ms and of 640 ns at 5 K. This permits several optical cycles to build spin polarisation and enhance the optical contrast up to 14. The coherent properties can be significantly improved using a non-isostructural Sn(4-fluoro-2-methylphenyl)4 host matrix [143]. Here, the ZFS is significantly increased, giving rise to clock transitions similar to those realised by the so-called "zero first-order Zeeman" (ZEFOZ) method, initially pioneered for lanthanide spin systems (see below). As a result, and are increased to 1.210.02 ms and 10.60.2 s at 5 K, respectively, while the pulsed contrast can be improved up to 65.
Work continues to explore alternative d-block elements where strong ligand-field interactions can enable optical excitation between spin manifolds and spin-dependent fluorescence [115, 118]. Candidate include Ni2+ [144, 145], Mo4+ [146], V3+ [147, 146]. Current challenges involve tuning the energy of the intermediate (e.g., singlet) state to provide the right conditions for spin polarisation, and identifying the appropriate ligand design to optimise photoluminescence.
2.3.3 f-block molecular spin centres
Lastly, there has been significant interest in f-block molecular systems. Here, we make a distinction between molecular systems where the ion is doped into a molecular lattice, and trapped ion systems where a single ion is levitated using electrostatic interactions to give rise to extremely long coherence times, but are not subject to qubit-quality improvements through molecular engineering [148, 149, 150].
Lanthanide spin centres benefit from highly shielded f-orbital electrons compared to d-orbital systems, leading to relatively long coherence times at low temperatures [151]. The core-like nature of f-orbitals also means, compared to d-orbital systems, that the ligand-field environment has a much weaker influence on state energies (a few THz), compared to electron-electron interactions (800 - 1600 nm) and spin-orbit coupling (10s THz) [152, 153] (see Figure 9a). Lanthanide electron spins can also exhibit strong hyperfine coupling (tens of MHz) with I=5/2 (Eu, Pr, 173Yb), I=1/2 (171Yb), I=7/2 (149Sm, 167Er) nuclei, leading to a diverse spin-optical addressibility schemes that can be further enhanced using applied magnetic fields. Moreover, strong spin-orbit coupling can lead to high magnetic anisotropy that can protect spin states from small magnetic fluctuations [154, 155]. Examples of optically active materials benefit from narrow and stable spin-dependent emission profiles at near-infrared frequencies, making them potentially compatible with conventional telecom optical fibres. Colour centres for which optical addressability has been established include Ce3+ [156], Eu3+- [157], Pr3+-[50], Er3+- [47], Yb3+- [158], and Sm3+:Y2SiO5 [159] (Figure 9c).
@@
In the solid-state, Y2SiO5 has been favoured as host matrix due to the ability to grow large crystals with excellent optical properties using the Czochralski method [160] (Figure 9b). However, as the only naturally occurring isotope, 89Y harbours an I=1/2 nuclear spin that ultimately limits decoherence times. To reduce the impact of the spin bath, various decoupling techniques have emerged [161, 162, 45]. Perhaps the most successful is the so-called "ZEFOZ" technique, whereby a magnetic field is applied such that the magnetic-field dependence of the spin-transition frequency is very close to zero [163] (Figure 9d). In this "clock transition" regime, spins are first-order insensitive to small fluctuations in local magnetic fields. Using this technique with Eu3+:Y2SiO5, Zhong et al., demonstrated it is possible to acquire >>100 ms at 2 K, where spin-phonon coupling is negligible. Remarkably, combined with dynamic decoupling methods was measured to be 37060 mins, reaching a critical milestone whereby the distance-dependent decoherence becomes less for spin-transport than it is during light-transport of quantum states [164].
@@
Interestingly, the larger magnetic moment of Pr3+ can give rise to "frozen core"-type behaviour, whereby local Y-spins become dephased from the bulk crystal and hence exhibit reduced dephasing influence on the Pr3+ spins [161]. Equall et al., measured homogenous field-dependent and crystal structure site-dependent linewidths between 2.5 and 0.85 kHz, and a corresponding as high as 377 s at 1.4 K [50]. Combined with the ZEFOZ method, the can reach 82 ms at 1.5 K [161] (see example Figure 9e) and with further dynamical decoupling up to 1 min can be achieved, approaching the population lifetime limit [165]. Coherence times can also be enhanced using a host matrix whereby the principal host ion (e.g., Y) has a more closely matched ionic radius to the dopant, leading to reduced crystallographic distortions. For example, in a Pr3+:La2(WO4)3 system a of 1587 ms has been measured at 4 K [51].
@@
Spin coherence times of Er3+:Y2SiO5 have so far been shorter than the best performing Eu3+ materials with measured at 1.30.01 seconds at 1.4 K, despite exhibiting the frozen core effect [47]. However, this was achieved without using the ZEFOZ method, and therefore, these times can likely be significantly extended. Er3+ has also been studied in a Y2O3 host with reaching 140 s at 1.8 K and with an applied field of 4 T [46]. Decoherence was dominated by phonon-driven dipole-dipole interactions and the nuclear spin bath at high fields, similar to Er3+:KTiOPO4 where was measured at 200 s under similar conditions [48].
@@
171Yb3+ holds a unique position amongst the lanthanide ions discussed so far due to its S=1/2 and I=1/2 electron spin and hyperfine structure, resulting in a simple 4-level system. Of the 171Yb3+-doped materials [53], 171Yb3+:Y2SiO5 appears to exhibit the longest spin coherence times. This material was initially studied by X-band EPR spectroscopy and presented with an electron of 5 seconds at 2.5 K with an applied field of 90 mT. quickly increases above 4 K due to Raman relaxation where becomes ultimately limited by [52]. At 2.5 K, was optimised at 1 T to 73 s and improved further by dynamical decoupling reaching 550 s. In the same experiments, the authors record nuclear and at 4.5 K of 4 and 0.35 ms, respectively. Using an optical approach, Ortu et al., improved the coherent properties of 171Yb3+:Y2SiO5 by employing the ZEFOZ method such that the electron remains above 100 s at 5.6 K and the nuclear extends to 1 ms [158]. Using a different approach, Welinski et al., demonstrated that coherence can also be extended by first polarising host nuclear spins through spin diffusion. At 2 K, the authors first excite 171Yb spins before allowing them to equilibrate over a few seconds through spectral diffusion over the inhomogeneous linewidth. The result is an effective "hole burning" in the absorption spectrum of 171Yb3+:Y2SiO5 and up to 90 nuclear spin polarisation, thereby effectively generating mK spin temperatures and improving the optical from 0.3 to 0.8 ms [168].
3 Optimising Performance of Spin-Qubit Materials